# Quantum (quadratic) gravity: replacing the massive tensor ghost with an inverted harmonic oscillator-like instability

### 07 Mar 2026 v1

@@ 7.2 Parity asymmetry of primordial power spectra
In the exact de Sitter spacetime, the Bunch–Davies vacuum can be defined by analytic continuation from the Euclidean four-sphere [7]. The regularity of the Euclidean correlator enforces periodicity in imaginary conformal time with period
$\beta=\frac{2\pi}{|H|}.$ (7.59)
@@
$G^{>}(t-i\beta,\mathbf{k})=G^{<}(t,\mathbf{k})\quad{\rm or}\quad G^{>}\!\left(\tau e^{i|H|\beta},\tau^{\prime};k\right)=G^{<}\!\left(\tau,\tau^{\prime};k\right)$ (7.63)
@@
$\tau=-\frac{1}{H}e^{-Ht}\quad\Rightarrow\quad t\to t-i\beta,\qquad\beta=\frac{2\pi}{H}\quad\Longleftrightarrow\quad\tau\to e^{2\pi i}\tau.$ (7.68)
@@
$G_{k}^{>}(t-i\beta,\tau^{\prime})=e^{-2\pi i\nu}\,G_{k}^{>}(\tau,\tau^{\prime}).$ (7.70)
@@
$G_{k}^{>}(t-i\beta,\tau^{\prime})=G_{k}^{<}(t,\tau^{\prime}).$ (7.71)
@@
$G_{k}^{>,\pm}(t-i\beta,\tau^{\prime})=e^{\mp 4\pi i\epsilon_{1}}\,G_{k}^{<,\pm}(t,\tau^{\prime})+\mathcal{O}(\epsilon_{1}^{2}).$ (7.76)
@@
$G_{k}^{>,\text{tot}}(t-i\beta)=\cos(4\pi\epsilon_{1})\,G_{k}^{<,\text{tot}}(t)+\mathcal{O}(\epsilon_{1}^{2}).$ (7.78)
@@ 8 Conclusions
To be more explicit, we showed that the spin–2 ghost of quadratic gravity can be replaced by an IHO-like (called "dual IHO") spin–2 degree of freedom rather than a pathological negative norm excitation (i.e, ghost). This occurs by a change in sign of the coefficient of Weyl square term in the quadratic gravity action (4.1). This approach replaces the conventional ghost instability by a healthy dynamical instability associated with saddle point dynamics, which is also prevalent in the contexts of the SM Higgs mechanism and quantum fields near gravitational horizons.
@@
Building on this insight, the direct-sum quantization framework we developed earlier brings here a consistent understanding of quantum IHO(-like) degrees of freedom. Within this formulation, the spin–2 transverse traceless sector of quadratic gravity admits a unitary quantization, restoring probabilistic consistency while preserving renormalizability. We demonstrated explicitly that scattering amplitudes satisfy the optical theorem. We clarified how imaginary parts do not arise through standard cutting identities (i.e., Cutkowsky rules) because of the spacelike support of the IHO-like modes. Our analysis further suggests that the massive spin–2 sector, when treated as a dual IHO system, effectively decouples from asymptotic observables while remaining dynamically relevant in the deep ultraviolet virtual quantum processes. This provides a concrete mechanism by which quadratic gravity can remain unitary without introducing additional propagating polarizations in physical graviton states. Embracing the finite action principle, unitary QQG might be free from cosmological and BH singularities. From a cosmological standpoint, the dual IHO dynamics offers a natural route toward singularity avoidance and a nonsingular origin of the Universe, supplying a driving instability that replaces classical divergences by quantum controlled evolution. Importantly, the IHO-like instability in quadratic gravity has a universal appeal. The same IHO structure (which is nothing but the Berry–Keating $H=xp$ Hamiltonian [29]), appears in the descriptions of near-horizon QFTCS, in tachyonic nature of $\mathbb{Z}_{2}$ symmetry breaking of the SM, in inflationary (quasi-de Sitter) perturbations. Interestingly Berry-Keating IHO is also widely discussed in other fields such as condensed matter physics. Thus, the recurrence of this structure across such disparate areas strongly suggests that the inverted harmonic oscillator is not an anomaly of higher-derivative gravity, but a universal dynamical template of unstable or horizon-dominated systems in fundamental physics.
@@
More broadly, our results highlight a unifying role played by IHOs across quantum field theory, gravitational horizons, and UV behavior of higher-derivative gravity. Therefore, rather than signaling inconsistency, inverted oscillator dynamics encodes horizon physics, tachyonic phases, and ultraviolet gravitational behavior within a single conceptual framework. This perspective opens a viable path toward a unitary, renormalizable theory of gravity and motivates further investigation into its observational consequences, including early Universe dynamics and tensor sector signatures. Interestingly, with dual IHO spin–2 field, the renormalization group structure of QQG is analogous to that of QED i.e., QQG remains perturbatively well defined and predictive across all physically relevant scales because the Landau pole (as per 1-loop beta function) in QQG appears parametrically far beyond the Planck regime, where new physics or nonperturbative effects are expected to intervene. The presence of the spin–2 IHO degree of freedom does not destabilize this structure; rather, it participates in the UV running while remaining absent from asymptotic spectra. On the observational side, the same sector leaves a controlled imprint in the early Universe. In particular, the tensor-to-scalar ratio receives a correction

### 25 Mar 2026 v2

@@ 7.2 Parity asymmetry of primordial power spectra
In the exact de Sitter spacetime, the Bunch–Davies vacuum can be defined by analytic continuation from the Euclidean four-sphere [7]. The regularity of the Euclidean correlator enforces periodicity in imaginary conformal time with period [^11] Note that the notation $\beta=1/T$ here refers to the inverse of the temperature and it should not be confused with the coefficient of Weyl square in quadratic gravity (4.1) that we discussed earlier.
$\beta_{T}=\frac{2\pi}{|H|}.$ (7.59)
@@
$G^{>}(t-i\beta_{T},\mathbf{k})=G^{<}(t,\mathbf{k})\quad{\rm or}\quad G^{>}\!\left(\tau e^{i|H|\beta_{T}},\tau^{\prime};k\right)=G^{<}\!\left(\tau,\tau^{\prime};k\right)$ (7.63)
@@
$\tau=-\frac{1}{H}e^{-Ht}\quad\Rightarrow\quad t\to t-i\beta,\qquad\beta_{T}=\frac{2\pi}{|H|}\quad\Longleftrightarrow\quad\tau\to e^{2\pi i}\tau.$ (7.68)
@@
$G_{k}^{>}(t-i\beta_{T},\tau^{\prime})=e^{-2\pi i\nu}\,G_{k}^{>}(\tau,\tau^{\prime}).$ (7.70)
@@
$G_{k}^{>}(t-i\beta_{T},\tau^{\prime})=G_{k}^{<}(t,\tau^{\prime}).$ (7.71)
@@
$G_{k}^{>,\pm}(t-i\beta_{T},\tau^{\prime})=e^{\mp 4\pi i\epsilon_{1}}\,G_{k}^{<,\pm}(t,\tau^{\prime})+\mathcal{O}(\epsilon_{1}^{2}).$ (7.76)
@@
$G_{k}^{>,\text{tot}}(t-i\beta_{T})=\cos(4\pi\epsilon_{1})\,G_{k}^{<,\text{tot}}(t)+\mathcal{O}(\epsilon_{1}^{2}).$ (7.78)
@@ 8 Conclusions
To be more explicit, we showed that the spin–2 ghost of quadratic gravity can be replaced by an IHO-like (called "dual-IHO") spin–2 degree of freedom rather than a pathological negative norm excitation (i.e, ghost). This occurs by a change in sign of the coefficient of Weyl square term in the quadratic gravity action (4.1). This approach replaces the conventional ghost instability by a healthy dynamical instability associated with saddle point dynamics, which is also prevalent in the contexts of the SM Higgs mechanism and quantum fields near gravitational horizons.
Building on this insight, the direct-sum quantization framework we developed earlier brings here a consistent understanding of quantum IHO(-like) degrees of freedom. Within this formulation, the spin–2 transverse traceless sector of quadratic gravity admits a unitary quantization, restoring probabilistic consistency while preserving renormalizability. We demonstrated explicitly that scattering amplitudes satisfy the optical theorem. We clarified how imaginary parts do not arise through standard cutting identities (i.e., Cutkowsky rules) because of the spacelike support of the IHO-like modes. Our analysis further suggests that the massive spin–2 sector, when treated as a dual-IHO system, effectively decouples from asymptotic observables while remaining dynamically relevant in the deep ultraviolet virtual quantum processes. This provides a concrete mechanism by which quadratic gravity can remain unitary without introducing additional propagating polarizations in physical graviton states. A particularly sharp result of our construction is the derivation of the dual-IHO propagator from first principles via the Källén–Lehmann representation. The spectral density $\rho$ vanishes identically for two independent and self-consistent reasons: first, the dual-IHO ground state is not square-integrable, so no normalizable vacuum exists from which to define the KL matrix elements; second, the propagator pole sits at spacelike momentum $k^{2}=\mu_{2}^{2}>0$, which lies strictly outside the timelike integration domain of the KL representation. Together, these enforce the principal value propagator as the unique physically consistent prescription, not as an external input, but as a theorem derived from the Hilbert space structure of the theory. This is the precise mechanism by which the dual-IHO spin–2 mode contributes to UV renormalization through real dispersive loop effects, while contributing zero imaginary part to any amplitude at any loop order, thereby leaving the optical theorem intact.
@@
Embracing the finite action principle, unitary QQG might be free from cosmological and BH singularities. From a cosmological standpoint, the dual-IHO dynamics offers a natural route toward singularity avoidance and a nonsingular origin of the Universe, supplying a driving instability that replaces classical divergences by quantum controlled evolution. Importantly, the IHO-like instability in quadratic gravity has a universal appeal. The same IHO structure (which is nothing but the Berry–Keating $H=xp$ Hamiltonian [29]), appears in the descriptions of near-horizon QFTCS, in tachyonic nature of $\mathbb{Z}_{2}$ symmetry breaking of the SM, in inflationary (quasi-de Sitter) perturbations. Interestingly Berry-Keating IHO is also widely discussed in other fields such as condensed matter physics. Thus, the recurrence of this structure across such disparate areas strongly suggests that the inverted harmonic oscillator is not an anomaly of higher-derivative gravity, but a universal dynamical template of unstable or horizon-dominated systems in fundamental physics.
@@
More broadly, our results highlight a unifying role played by IHOs across quantum field theory, gravitational horizons, and UV behavior of higher-derivative gravity. Therefore, rather than signaling inconsistency, inverted oscillator dynamics encodes horizon physics, tachyonic phases, and ultraviolet gravitational behavior within a single conceptual framework. This perspective opens a viable path toward a unitary, renormalizable theory of gravity and motivates further investigation into its observational consequences, including early Universe dynamics and tensor sector signatures. Interestingly, with dual-IHO spin–2 field, the renormalization group structure of QQG is analogous to that of QED i.e., QQG remains perturbatively well defined and predictive across all physically relevant scales because the Landau pole (as per 1-loop beta function) in QQG appears parametrically far beyond the Planck regime, where new physics or nonperturbative effects are expected to intervene. The presence of the spin–2 IHO degree of freedom does not destabilize this structure; rather, it participates in the UV running while remaining absent from asymptotic spectra. On the observational side, the same sector leaves a controlled imprint in the early Universe. In particular, the tensor-to-scalar ratio receives a correction